OpenStudy (anonymous):

Jim Ross is an automobile salesman who receives a salary of $500 per week plus a commission of 3 percent on all sales. During a 4-week period he sold $42,000 worth of cars. What were Jim's average earnings?

OpenStudy (anonymous):

salary of 4 weeks = 4 x 500 = 2000 commission of 4 weeks = 0.03 x 42000 = 1260 so total earning for 4 weeks = 2000 + 1260 = 3260 avg earning per week = 3260/4 = 815